# The Fragile Nature of Road Transportation Systems

### v2

@@ 4.1 Demand disruption
$=-\frac{e^{-a_{2}t}(m_{c}-q)}{a_{2}^{2}}e^{a_{2}t_{c}}+\frac{b_{2}-q}{a_{2}}t_{c}+\frac{m_{c}-q}{a_{2}^{2}}-\frac{b_{2}-q}{a_{2}}t$ (39)
The derivatives on the less congested branch are:
$\frac{dTTS_{2}}{dn^{\prime}}$ $=-\frac{(m_{c}-q)^{1-\frac{a_{2}}{a_{1}}}e^{-a_{2}t}}{a_{2}}(a_{1}n^{\prime}+b_{1}-q)^{\frac{a_{2}}{a_{1}}-1}+\frac{b_{2}-q}{a_{2}}(a_{1}n^{\prime}+b_{1}-q)^{-1}$ (40)
$\frac{d^{2}TTS_{2}}{dn^{\prime 2}}$ $=-\left(\frac{(a_{2}-a_{1})(m_{c}-q)^{1-\frac{a_{2}}{a_{1}}}e^{-a_{2}t}}{a_{2}}(a_{1}n^{\prime}+b_{1}-q)^{\frac{a_{2}}{a_{1}}}+\frac{a_{1}(b_{2}-q)}{a_{2}}\right)(a_{1}n^{\prime}+b_{1}-q)^{-2}$ (41)
The second derivative of the whole process can be written as:
$\frac{d^{2}TTS}{dn^{\prime 2}}$ $=\frac{d^{2}TTS_{1}}{dn^{\prime 2}}+\frac{d^{2}TTS_{2}}{dn^{\prime 2}}$ (42)
$=\left(b_{1}-q-\frac{e^{-a_{2}t}}{a_{2}}(a_{2}-a_{1})(m_{c}-q)^{1-\frac{a_{2}}{a_{1}}}(m_{0}-q)^{\frac{a_{2}}{a_{1}}}-\frac{a_{1}(b_{2}-q)}{a_{2}}\right)(m_{0}-q)^{-2}$ (43)
@@
As per Assumption 3, we have $m_{0}-q>0$, so if a transportation system is to be fragile, $d^{2}TTS/dn^{\prime 2}$ should also be positive, and the following equation has to be true:
$b_{1}-q-\frac{e^{-a_{2}t}}{a_{2}}(a_{2}-a_{1})(m_{c}-q)^{1-\frac{a_{2}}{a_{1}}}(m_{0}-q)^{\frac{a_{2}}{a_{1}}}-\frac{a_{1}(b_{2}-q)}{a_{2}}>0$ (44)
@@
Since $t>t_{c}$ and $a_{2}>a_{1}$, regardless of whether $a_{2}$ is positive or negative, the following relationship always holds:
$-\frac{e^{-a_{2}t}}{a_{2}}>-\frac{e^{-a_{2}t_{c}}}{a_{2}}$ (45)
@@
As the following three terms, $a_{2}-a_{1}$, $(m_{c}-q)^{1-\frac{a_{2}}{a_{1}}}$, and $(m_{0}-q)^{\frac{a_{2}}{a_{1}}}$ are all positive, the following relationship is true:
$b_{1}-q-\frac{e^{-a_{2}t}}{a_{2}}(a_{2}-a_{1})(m_{c}-q)^{1-\frac{a_{2}}{a_{1}}}(m_{0}-q)^{\frac{a_{2}}{a_{1}}}-\frac{a_{1}(b_{2}-q)}{a_{2}}>$ (46)
$b_{1}-q-\frac{e^{-a_{2}t_{c}}}{a_{2}}(a_{2}-a_{1})(m_{c}-q)^{1-\frac{a_{2}}{a_{1}}}(m_{0}-q)^{\frac{a_{2}}{a_{1}}}-\frac{a_{1}(b_{2}-q)}{a_{2}}$ (47)
Here we substitute $t_{c}$ in Eq. 47 with Eq. 31 and we get:
$b_{1}-q-\frac{(a_{2}-a_{1})(m_{c}-q)}{a_{2}}-\frac{a_{1}(b_{2}-q)}{a_{2}}$ $=$ (48)
$a_{1}\left(\frac{b_{1}-m_{c}}{a_{1}}-\frac{b_{2}-m_{c}}{a_{2}}\right)=a_{1}(n_{c}-n_{c})$ $=0$ (49)
Hence, we have:
$b_{1}-q-\frac{e^{-a_{2}t}}{a_{2}}(a_{2}-a_{1})(m_{c}-q)^{1-\frac{a_{2}}{a_{1}}}(m_{0}-q)^{\frac{a_{2}}{a_{1}}}-\frac{a_{1}(b_{2}-q)}{a_{2}}>0$ (50)
@@ 4.2 Supply disruption
While a positive second derivative of ATS or TTS on traffic demand signifies the fragility of the road transportation network to demand disruptions, establishing a positive second derivative of time spent concerning the magnitude of FD/MFD disruption would demonstrate the fragility from the perspective of supply disruptions. Here we use a supply disruption magnitude coefficient $r$ and the disrupted FD/MFD is expressed as $(1-r)G(n)$. Although real-world MFD may be decreased in various shapes, we use this simple approach as applied in Ambühl et al. (2020) when studying the uncertainties of MFDs. The physical meaning of $(1-r)G(n)$ relates to the decrease of the free-flow speed due to, e.g., snowy weather and icy roads, with the maximal density of the network remaining unchanged. The traffic demand at equilibrium before the MFD disruption is $q=G(k_{0})$, or $q=G(n_{0})$. After the supply disruption, as per Assumption 3, the supply disruption magnitude coefficient $r\in[0,1)$ is not significantly large so the traffic demand remains below the maximal capacity on the disrupted MFD profile, and the new equilibrium point is $q=(1-r)G(k^{\prime}(r))$, or $q=(1-r)G(n^{\prime}(r))$. It should be noted that, unlike the study of demand disruption, when studying supply disruptions, $k^{\prime}(r)$ is a dependent variable on $r$.
@@
Road transportation systems are fragile with the onset of supply disruptions on the microscopic level.
@@
First, we study the fragile properties of the road transportation systems under the onset of supply disruptions based on the microscopic model Greenshields FD. Here, the base demand $q$ is constant before and after the onset of the supply disruption, and can be used to build the relationship between the two stable states before and after the supply disruption.
$q=G(k_{0})=(1-r)G(k^{\prime}(r))=(1-r)(ak^{\prime}(r)^{2}+bk^{\prime}(r))$ (51)
So, the traffic density of the new equilibrium point after the MFD disruption would be:
$k^{\prime}(r)$ $=\frac{\sqrt{b^{2}+\frac{4aq}{1-r}}-b}{2a}$ (52)
@@
$ATS$ $=\frac{L}{v(k^{\prime})}=\frac{Lk^{\prime}(r)}{q}=\frac{L}{2aq}((b^{2}+\frac{4aq}{1-r})^{\frac{1}{2}}-b)$ (53)
$\frac{dATS}{dr}$ $=L(b^{2}(1-r)+4aq)^{-\frac{1}{2}}(1-r)^{-\frac{3}{2}}$ (54)
$\frac{d^{2}ATS}{dr^{2}}$ $=\frac{b^{2}L}{2}(b^{2}(1-r)+4aq)^{-\frac{3}{2}}(1-r)^{-\frac{3}{2}}+\frac{3L}{2}(b^{2}(1-r)+4aq)^{-\frac{1}{2}}(1-r)^{-\frac{5}{2}}$ (55)
As $k^{\prime}$ has a physical meaning of disruption density, so it should have real roots with $b^{2}+4aq/1-r$ being positive. And since $01-r>$ 0, therefore, $d^{2}ATS/dr^{2}$ is always positive and indicates the fragility property of transportation systems when faced with supply disruption on the microscopic level. Likewise to the proof of demand disruption, the Daganzo FD is a special case of the MoC, so we directly prove the fragility on the macroscopic level with MoC illustrated in the following section.
@@
Road transportation systems are fragile with the onset of supply disruptions on the macroscopic level.
@@
As the proof of fragility under supply disruption with cubic polynomial MFDs involves solving the roots for cubic equations, for simplicity reasons, we prove only with Daganzo MoC. Since the traffic demand $q$ is constant, the traffic density of the new stable state $k^{\prime}(r)$ would be:
$q$ $=(1-r)(uk^{\prime}(r)+c)$ (56)
$k^{\prime}(r)$ $=\frac{\frac{q}{1-r}-c}{u}$ (57)
@@
$ATS$ $=\frac{Lk^{\prime}(r)}{q}=\frac{L}{qu}(\frac{q}{1-r}-c)$ (58)
$\frac{dATS}{dr}$ $=\frac{L}{u}(1-r)^{-2}$ (59)
$\frac{d^{2}ATS}{dr^{2}}$ $=\frac{2L}{u}(1-r)^{-3}$ (60)
@@
To study the possible fragile properties of road transportation networks regarding the recovery process from supply disruptions, we need to combine the conclusions from Proposition 3 and Proposition 5. In Proposition 3, we’ve proven the recovery process to be fragile when the traffic state shifts from the more congested branch to the less congested branch for any two consecutive cuts on the Daganzo MFD, or to be neither fragile nor antifragile when it stays only on one single branch. Therefore, it can be mathematically summarized as $d^{2}TTS/dn^{\prime 2}\geq 0$. Following Assumption 2, $a_{1}>0$ holds as the branch is below the critical density, we can easily prove the first derivative $dTTS/dn^{\prime}$ to be non-negative as well with Eq. 34 and Eq. 40 through the same procedure as proving the second derivative to be positive. And for the proof of recovery from supply disruptions, as shown in Fig. 8, likewise to Proposition 5, the relationship between the original and new equilibrium points before and after MFD disruptions can be expressed as:
@@
$q$ $=un_{0}+c=(1-r)(un^{\prime}+c)$ (61)
$n^{\prime}(r)$ $=\frac{un_{0}+c}{u(1-r)}-c/u$ (62)
@@
$\frac{dn^{\prime}}{dr}$ $=\frac{un_{0}+c}{u}(1-r)^{-2}$ (63)
$\frac{d^{2}n^{\prime}}{dr^{2}}$ $=\frac{2(un_{0}+c)}{u}(1-r)^{-3}$ (64)
@@
Since $u$ and $un_{0}+c$ are both positive, the derivatives of $n^{\prime}$ over $r$ are positive as well. Additionally, it can be easily proven that when considering the transition from a less congested to a more congested branch, the same conclusion also holds. As $TTS$ is a function of $n^{\prime}$ and $n^{\prime}$ is again a function of $r$, by applying the chain rule, we can get the second derivative of $TTS$ over $r$ as:
$\frac{d^{2}TTS}{dr^{2}}$ $=\frac{d}{dr}\left(\frac{dTTS}{dn^{\prime}}\cdot\frac{dn^{\prime}}{dr}\right)$ (65)
$=\frac{d^{2}TTS}{dn^{\prime 2}_{0}}\cdot\left(\frac{dn^{\prime}}{dr}\right)^{2}+\frac{dTTS}{dn^{\prime}}\cdot\frac{d^{2}n^{\prime}}{dr^{2}}$ (66)

### v4

@@ 4.1 Demand disruption
Therefore, with the disruption accumulation $n^{\prime}$, and when the traffic states are assumed to be on the same cut $i$. After any time $t$, the vehicle accumulation $n$ would be:
$n$ $=\frac{a_{i}n^{\prime}+b_{i}-m_{0}}{a_{i}}e^{-a_{i}t}-\frac{b_{i}-m_{0}}{a_{i}}$ (14)
The TTS on this cut $i$ for a given period $t$ can be calculated as:
[…]
$=-\frac{a_{i}n^{\prime}+b_{i}-m_{0}}{a_{i}^{2}}e^{-a_{i}t}-\frac{b_{i}-m_{0}}{a_{i}}t+\frac{a_{i}n^{\prime}+b_{i}-m_{0}}{a_{i}^{2}}$ (15)
Now we calculate the derivatives of TTS assuming $t<t_{\tilde{c},i}$.
$\frac{dTTS}{dn^{\prime}}$ $=\frac{1}{a_{i}}-\frac{e^{-a_{i}t}}{a_{i}}$ (16a)
$\frac{d^{2}TTS}{dn^{\prime 2}}$ $=0$ (16b)
The second derivative of TTS is 00, indicating that when the traffic states move only along a single cut $i$, it shows neither fragility nor antifragility. On the other hand, when the traffic state goes over the critical vehicle accumulation $n_{\tilde{c},i}$, and since the MoC is a piecewise function, we calculate the TTS and the related derivatives separately on each concerned cut $y,y+1,\dots,z$, denoted as $TTS_{y},TTS_{y+1},\dots,TTS_{z}$. Since the critical time $t_{\tilde{c},i}$ is yet unknown, we need to determine $t_{\tilde{c},i}$ first based on Eq. 13b.
$t_{\tilde{c},y}$ $=-\frac{1}{a_{y}}\ln{\left(\frac{a_{y}n_{\tilde{c},y}+b_{y}-m_{0}}{a_{y}n^{\prime}+b_{y}-m_{0}}\right)}$ (17a)
$t_{\tilde{c},y+1}-t_{\tilde{c},y}$ $=-\frac{1}{a_{y+1}}\ln{\left(\frac{a_{y+1}n_{\tilde{c},{y+1}}+b_{y+1}-m_{0}}{a_{y+1}n_{\tilde{c},y}+b_{y+1}-m_{0}}\right)}$ (17b)
[…]
$t_{\tilde{c},z-1}-t_{\tilde{c},z-2}$ $=-\frac{1}{a_{z-1}}\ln{\left(\frac{a_{z-1}n_{\tilde{c},{z-1}}+b_{z-1}-m_{0}}{a_{z-1}n_{\tilde{c},z-2}+b_{z-1}-m_{0}}\right)}$ (17c)
Therefore, the time $t_{\tilde{c},z-1}$ to reach the last critical point $n_{\tilde{c},z-1}$ is:
$t_{\tilde{c},z-1}$ $=-\frac{1}{a_{y}}\ln{\left(\frac{a_{y}n_{\tilde{c},y}+b_{y}-m_{0}}{a_{y}n^{\prime}+b_{y}-m_{0}}\right)}-\sum_{i=y+1}^{z-1}\frac{1}{a_{i}}\ln{\left(\frac{a_{i}n_{\tilde{c},{i}}+b_{i}-m_{0}}{a_{i}n_{\tilde{c},i-1}+b_{i}-m_{0}}\right)}$ (18)
As $a_{y}n_{\tilde{c},y}+b_{y}$ is equal to $m_{\tilde{c},y}$, we can rewrite the above Eq. 17a as:
$t_{\tilde{c},y}$ $=-\frac{1}{a_{y}}\ln{\left(\frac{m_{\tilde{c},y}-m_{0}}{a_{y}n^{\prime}+b_{y}-m_{0}}\right)}$ (19)
A generalized form between any two consecutive critical points $n_{\tilde{c},i-1}$ and $n_{\tilde{c},i}$ will be:
$t_{\tilde{c},i}-t_{\tilde{c},i-1}$ $=-\frac{1}{a_{i}}\ln{\left(\frac{m_{\tilde{c},{i}}-m_{0}}{m_{\tilde{c},i-1}-m_{0}}\right)}$ (20)
Similarly, the above Eq. 18 can also be rewritten as:
$t_{\tilde{c},z-1}$ $=-\frac{1}{a_{y}}\ln{\left(\frac{m_{\tilde{c},y}-m_{0}}{a_{y}n^{\prime}+b_{y}-m_{0}}\right)}-\sum_{i=y+1}^{z-1}\frac{1}{a_{i}}\ln{\left(\frac{m_{\tilde{c},i}-m_{0}}{m_{\tilde{c},i-1}-m_{0}}\right)}$ (21)
We substitute $t$ in Eq. 15 with $t_{\tilde{c},y}$ in Eq. 19, and the $TTS_{y}$ for the more congested cut will be:
[…]
$=-\frac{m_{\tilde{c},y}-m_{0}}{a_{y}^{2}}+\frac{b_{y}-m_{0}}{a_{y}^{2}}\ln{\left(\frac{m_{\tilde{c},y}-m_{0}}{a_{y}n^{\prime}+b_{y}-m_{0}}\right)}+\frac{a_{y}n^{\prime}+b_{y}-m_{0}}{a_{y}^{2}}$ (22)
The derivatives for $TTS_{y}$ are:
$\frac{dTTS_{y}}{dn^{\prime}}$ $=-\frac{b_{y}-m_{0}}{a_{y}}(a_{y}n^{\prime}+b_{y}-m_{0})^{-1}+\frac{1}{a_{y}}$ (23a)
$\frac{d^{2}TTS_{y}}{dn^{\prime 2}}$ $=(b_{y}-m_{0})(a_{y}n^{\prime}+b_{y}-m_{0})^{-2}$ (23b)
Now we consider the cuts between the most congested cut $y$ and the least congested cut $z$. Combining Eq. 13c and Eq. 15, the $TTS_{i}$ on any cut $y<i<z$ is:
[…]
$=-\frac{e^{-a_{i}(t_{\tilde{c},i}-t_{\tilde{c},i-1})}(m_{\tilde{c},i-1}-m_{0})}{a_{i}^{2}}+\frac{(m_{\tilde{c},i-1}-m_{0})}{a_{i}^{2}}-\frac{b_{i}-m_{0}}{a_{i}}(t_{\tilde{c},i}-t_{\tilde{c},i-1})$ (24)
Then $t_{\tilde{c},i}-t_{\tilde{c},i-1}$ can be substituted with Eq. 20:
$TTS_{i}=-\frac{m_{\tilde{c},i}-m_{\tilde{c},i-1}}{a_{i}^{2}}+\frac{b_{i}-m_{0}}{a_{i}^{2}}\ln{\left(\frac{m_{\tilde{c},{i}}-m_{0}}{m_{\tilde{c},i-1}-m_{0}}\right)}$ (25)
As Eq. 25 is not dependent on $n^{\prime}$. Therefore, the second derivative of $TTS_{i}$ over $n^{\prime}$ with $y<i<z$ are:
$\frac{dTTS_{i}}{dn^{\prime}}$ $=0$ (26a)
$\frac{d^{2}TTS_{i}}{dn^{\prime 2}}$ $=0$ (26b)
The final piece of the puzzle is the least congested cut $z$, and likewise to Eq. 24, the $TTS_{z}$ is:
$TTS_{z}=-\frac{e^{-a_{z}(t-t_{\tilde{c},z-1})}(m_{\tilde{c},z-1}-m_{0})}{a_{z}^{2}}+\frac{(m_{\tilde{c},z-1}-m_{0})}{a_{z}^{2}}-\frac{b_{z}-m_{0}}{a_{z}}(t-t_{\tilde{c},z-1})$ (27)
We substitute $t_{\tilde{c},z-1}$ with Eq. 21:
$TTS_{z}=$ $-\frac{e^{-a_{z}t}}{a_{z}^{2}}(m_{\tilde{c},z-1}-m_{0})\left(\frac{m_{\tilde{c},y}-m_{0}}{a_{y}n^{\prime}+b_{y}-m_{0}}\right)^{-\frac{a_{z}}{a_{y}}}\prod_{i=y+1}^{z-1}\left(\frac{m_{\tilde{c},i}-m_{0}}{m_{\tilde{c},i-1}-m_{0}}\right)^{-\frac{a_{z}}{a_{i}}}+\frac{(m_{\tilde{c},z-1}-m_{0})}{a_{z}^{2}}$
$-\frac{b_{z}-m_{0}}{a_{z}}\left(t+\frac{1}{a_{y}}\ln{\left(\frac{a_{y}n_{\tilde{c},y}+b_{y}-m_{0}}{a_{y}n^{\prime}+b_{y}-m_{0}}\right)}+\sum_{i=y+1}^{z-1}\frac{1}{a_{i}}\ln{\left(\frac{a_{i}n_{\tilde{c},{i}}+b_{i}-m_{0}}{a_{i}n_{\tilde{c},i-1}+b_{i}-m_{0}}\right)}\right)$ (28)
A constant $P$ is introduced for conciseness since all the parameters involved below are constants:
$P=(m_{\tilde{c},y}-m_{0})^{-\frac{a_{z}}{a_{y}}}\prod_{i=y+1}^{z-1}\left(\frac{m_{\tilde{c},i}-m_{0}}{m_{\tilde{c},i-1}-m_{0}}\right)^{-\frac{a_{z}}{a_{i}}}$ (29)
The derivatives for the least congested cut $z$ are:
$\frac{dTTS_{z}}{dn^{\prime}}$ $=-\frac{e^{-a_{z}t}P}{a_{z}}(m_{\tilde{c},z-1}-m_{0})(a_{y}n^{\prime}+b_{y}-m_{0})^{\frac{a_{z}}{a_{y}}-1}+\frac{b_{z}-m_{0}}{a_{z}}(a_{y}n^{\prime}+b_{y}-m_{0})^{-1}$ (30a)
$\frac{d^{2}TTS_{z}}{dn^{\prime 2}}$ $=-\left(\frac{e^{-a_{z}t}P}{a_{z}}(a_{z}-a_{y})(m_{\tilde{c},z-1}-m_{0})(a_{y}n^{\prime}+b_{y}-m_{0})^{\frac{a_{z}}{a_{y}}}+\frac{a_{y}(b_{z}-m_{0})}{a_{z}}\right)(a_{y}n^{\prime}+b_{y}-m_{0})^{-2}$ (30b)
The second derivative of the whole process is the sum of the second derivatives for the sub-process on each cut, as calculated in Eq. 23b, Eq. 26b, and Eq. 30b:
[…]
$=\left(b_{y}-m_{0}-\frac{e^{-a_{z}t}P}{a_{z}}(a_{z}-a_{y})(m_{\tilde{c},z-1}-m_{0})(m^{\prime}-m_{0})^{\frac{a_{z}}{a_{y}}}-\frac{a_{y}(b_{z}-m_{0})}{a_{z}}\right)(m^{\prime}-m_{0})^{-2}$ (31)
@@
As per Assumption 3, it is assumed that $m^{\prime}-m_{0}>0$, so if a transportation system is to be fragile, $\frac{d^{2}TTS}{dn^{\prime 2}}$ should be positive, and the following equation has to be proven to be true:
$b_{y}-m_{0}-\frac{e^{-a_{z}t}P}{a_{z}}(a_{z}-a_{y})(m_{\tilde{c},z-1}-m_{0})(m^{\prime}-m_{0})^{\frac{a_{z}}{a_{y}}}-\frac{a_{y}(b_{z}-m_{0})}{a_{z}}>0$ (32)
@@
Since $t>t_{\tilde{c},z-1}$, regardless of whether $a_{z}$ is positive or negative, the below relationship always holds:
$-\frac{e^{-a_{z}t}}{a_{z}}>-\frac{e^{-a_{z}t_{\tilde{c},z-1}}}{a_{z}}$ (33)
@@
As all of the following terms: $P$, $a_{z}-a_{y}$, $m_{\tilde{c},z-1}-m_{0}$, and $(m^{\prime}-m_{0})^{\frac{a_{z}}{a_{y}}}$, are all positive, the following relationship is true:
$b_{y}-m_{0}-\frac{e^{-a_{z}t}P}{a_{z}}(a_{z}-a_{y})(m_{\tilde{c},z-1}-m_{0})(m^{\prime}-m_{0})^{\frac{a_{z}}{a_{y}}}-\frac{a_{y}(b_{z}-m_{0})}{a_{z}}>$
$b_{y}-m_{0}-\frac{e^{-a_{z}t_{\tilde{c},z-1}}P}{a_{z}}(a_{z}-a_{y})(m_{\tilde{c},z-1}-m_{0})(m^{\prime}-m_{0})^{\frac{a_{z}}{a_{y}}}-\frac{a_{y}(b_{z}-m_{0})}{a_{z}}$ (34)
We substitute $t_{\tilde{c},z-1}$ in Eq. 34 with Eq. 21 and constant $P$ with Eq. 29 and here goes:
$b_{y}-m_{0}-\frac{(a_{z}-a_{y})(m_{\tilde{c},z-1}-m_{0})}{a_{z}}-\frac{a_{y}(b_{z}-m_{0})}{a_{z}}=b_{y}-m_{\tilde{c},z-1}-\frac{a_{y}}{a_{z}}(b_{z}-m_{\tilde{c},z-1})$ (35)
$m_{\tilde{c},z-1}$ can be substituted by $a_{z}n_{\tilde{c},z-1}+b_{z}$, so it further becomes:
$b_{y}-a_{z}n_{\tilde{c},z-1}-b_{z}-\frac{a_{y}}{a_{z}}(b_{z}-a_{z}n_{\tilde{c},z-1}-b_{z})=b_{y}-b_{z}+(a_{y}-a_{z})n_{\tilde{c},z-1}$ (36)
Since $z>y$, so both $b_{y}>b_{z}$ and $a_{y}<a_{z}$ hold true. Also, if cut $y$ and cut $z$ are consecutive, i.e., $z=y+1$, the virtual intersection point exists physically $n_{\check{c}}=n_{\tilde{c},z-1}$. Otherwise, when there is any other cut in between, then $n_{\check{c}}>n_{\tilde{c},z-1}$ holds. Therefore, it can be formulated as $n_{\check{c}}\geq n_{\tilde{c},z-1}$ and we define a non-negative value $\Delta n=n_{\check{c}}-n_{\tilde{c},z-1}$ so that Eq. 36 can be rewritten as:
$b_{y}-b_{z}+(a_{y}-a_{z})(n_{\check{c}}-\Delta n)$ $=(a_{y}n_{\check{c}}+b_{y})-(a_{z}n_{\check{c}}+b_{z})-(a_{y}-a_{z})\Delta n$
$=m_{\check{c}}-m_{\check{c}}-(a_{y}-a_{z})\Delta n\geq 0$ (37)
Having demonstrated its lower bound to be non-negative, it is certain that:
$\frac{d^{2}TTS}{dn^{\prime 2}}$ $=\left(b_{y}-m_{0}-\frac{e^{-a_{z}t}P}{a_{z}}(a_{z}-a_{y})(m_{\tilde{c},z-1}-m_{0})(m^{\prime}-m_{0})^{\frac{a_{z}}{a_{y}}}-\frac{a_{y}(b_{z}-m_{0})}{a_{z}}\right)(m^{\prime}-m_{0})^{-2}>0$ (38)
@@ 4.2 Supply disruption
While a positive second derivative of ATS or TTS on traffic demand signifies the fragility of the road transportation network to demand disruptions, establishing a positive second derivative of time spent concerning the magnitude of FD/MFD disruption would demonstrate the fragility from the perspective of supply disruptions. Here we use a supply disruption magnitude coefficient $r$ and the disrupted FD/MFD is expressed as $(1-r)G(n)$. Although real-world MFD may be decreased in various shapes, we use this simple approach as applied in Ambühl et al. (2020) when studying the uncertainties of MFDs. The physical meaning of $(1-r)G(n)$ relates to the decrease of the free-flow speed due to, e.g., snowy weather and icy roads, with the maximal density of the network remaining unchanged. The traffic demand at equilibrium before the MFD disruption is $q_{0}=G(k_{0})$, or $m_{0}=M(n_{0})$. After the supply disruption, as per Assumption 3, the supply disruption magnitude coefficient $r\in[0,1)$ is not significantly large so the traffic demand remains below the maximal capacity on the disrupted MFD profile, and the new equilibrium point is $q_{0}=(1-r)G(k^{\prime}(r))$, or $m_{0}=(1-r)M(n^{\prime}(r))$. It should be noted that, unlike the study of demand disruption, when studying supply disruptions, $k^{\prime}(r)$ is a dependent variable on $r$.
@@
Road transportation systems are fragile with the onset of supply disruptions at the link level.
@@
First, we study the fragile properties of the road transportation systems under the onset of supply disruptions based on the link-level Greenshields FD. Here, the base demand $q_{0}$ is constant before and after the onset of the supply disruption, and can be used to build the relationship between the two stable states before and after the supply disruption.
$q_{0}=G(k_{0})=(1-r)G(k^{\prime}(r))=(1-r)(\alpha_{2}k^{\prime}(r)^{2}+\alpha_{1}k^{\prime}(r))$ (39)
So, the traffic density of the new equilibrium point after the supply disruption would be:
$k^{\prime}(r)$ $=\frac{\sqrt{\alpha_{1}^{2}+\frac{4\alpha_{2}q_{0}}{1-r}}-\alpha_{1}}{2\alpha_{2}}$ (40)
@@
$ATS$ $=\frac{l}{v(k^{\prime})}=\frac{lk^{\prime}(r)}{q_{0}}=\frac{l}{2\alpha_{2}q_{0}}((\alpha_{1}^{2}+\frac{4\alpha_{2}q_{0}}{1-r})^{\frac{1}{2}}-\alpha_{1})$ (41a)
$\frac{dATS}{dr}$ $=l(\alpha_{1}^{2}(1-r)+4\alpha_{2}q_{0})^{-\frac{1}{2}}(1-r)^{-\frac{3}{2}}$ (41b)
$\frac{d^{2}ATS}{dr^{2}}$ $=\frac{\alpha_{1}^{2}l}{2}(\alpha_{1}^{2}(1-r)+4\alpha_{2}q_{0})^{-\frac{3}{2}}(1-r)^{-\frac{3}{2}}+\frac{3l}{2}(\alpha_{1}^{2}(1-r)+4\alpha_{2}q_{0})^{-\frac{1}{2}}(1-r)^{-\frac{5}{2}}$ (41c)
As $k^{\prime}$ has a physical meaning of disruption density, so it should have real roots with $\alpha_{1}^{2}+\frac{4\alpha_{2}q_{0}}{1-r}$ being positive. And since $01-r>$ 0, therefore, $\frac{d^{2}ATS}{dr^{2}}$ is always positive and indicates the fragility property of transportation systems when faced with supply disruption at the link level. Likewise to the proof of demand disruption, the Daganzo FD is a special case of the MoC, so we directly prove the fragility at the network level with MoC illustrated in the following section.
@@
Road transportation systems are fragile with the onset of supply disruptions at the network level.
@@
As the proof of fragility under supply disruption with cubic polynomial MFDs involves solving the roots for cubic equations, for simplicity reasons, we prove only with Daganzo MoC. Since the traffic demand $q_{0}$ is constant, the traffic density of the new stable state $k^{\prime}(r)$ would be:
$q_{0}$ $=(1-r)(u_{i}k^{\prime}(r)+c_{i})$ (42a)
$k^{\prime}(r)$ $=\frac{\frac{q_{0}}{1-r}-c_{i}}{u_{i}}$ (42b)
@@
$ATS$ $=\frac{lk^{\prime}(r)}{q_{0}}=\frac{l}{q_{0}u_{i}}\left(\frac{q_{0}}{1-r}-c_{i}\right)$ (43a)
$\frac{dATS}{dr}$ $=\frac{l}{u_{i}}(1-r)^{-2}$ (43b)
$\frac{d^{2}ATS}{dr^{2}}$ $=\frac{2l}{u_{i}}(1-r)^{-3}$ (43c)
@@
To study the possible fragile properties of road transportation networks regarding the recovery process from supply disruptions, we need to combine the conclusions from Proposition 3 and Proposition 5. In Proposition 3, we’ve proven the recovery process to be fragile when the traffic state shifts from a more congested cut to a less congested cut on the Daganzo MFD, or to be neither fragile nor antifragile when it stays only on one single cut. Therefore, it can be mathematically summarized as $\frac{d^{2}TTS}{dn^{\prime 2}}\geq 0$. Following Assumption 2, $a_{i}>0$ holds as the cut is below the critical density, we can easily prove the first derivative $\frac{dTTS}{dn^{\prime}}$ to be non-negative as well with Eq. 23a, Eq. 26a, and Eq. 30a through the same procedure as proving the second derivative to be positive. And for the proof of recovery from supply disruptions, as shown in Fig. 8, likewise to Proposition 5, the relationship between the original and new equilibrium points before and after MFD disruptions can be expressed as:
@@
$m_{0}$ $=u_{i}n_{0}+c_{i}=(1-r)(u_{i}n^{\prime}+c_{i})$ (44a)
$\Longrightarrow\hskip 16.0ptn^{\prime}(r)$ $=\frac{u_{i}n_{0}+c_{i}}{u_{i}(1-r)}-c_{i}/u_{i}$ (44b)
@@
$\frac{dn^{\prime}}{dr}$ $=\frac{u_{i}n_{0}+c_{i}}{u_{i}}(1-r)^{-2}$ (45a)
$\frac{d^{2}n^{\prime}}{dr^{2}}$ $=\frac{2(u_{i}n_{0}+c_{i})}{u_{i}}(1-r)^{-3}$ (45b)
@@
Since $u_{i}$ and $u_{i}n_{0}+c_{i}$ are both positive, the derivatives of $n^{\prime}$ over $r$ are positive as well. Additionally, it can be easily proven that when considering the transition from a less congested to a more congested cut, the same conclusion also holds. As $TTS$ is a function of $n^{\prime}$ and $n^{\prime}$ is again a function of $r$, by applying the chain rule, we can get the second derivative of $TTS$ over $r$ as:
$\frac{d^{2}TTS}{dr^{2}}$ $=\frac{d}{dr}\left(\frac{dTTS}{dn^{\prime}}\cdot\frac{dn^{\prime}}{dr}\right)$
$=\frac{d^{2}TTS}{dn^{\prime 2}}\cdot\left(\frac{dn^{\prime}}{dr}\right)^{2}+\frac{dTTS}{dn^{\prime}}\cdot\frac{d^{2}n^{\prime}}{dr^{2}}$ (46)